# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v3

@@ 3 Theory
This section establishes a rigorous theoretical foundation for integrating SO(3)-equivariance with non-linear expressiveness in the deep learning paradigm. First, we propose a theoretical construct to derive SO(3)-invariant quantities from the SO(3)-equivariant targets, aiming at facilitating the learning of high-quality SO(3)-invariant features. Second, we propose a gradient-based mechanism encoding features maintaining theoretically SO(3)-equivariance while possessing the non-linear expressive capability induced from SO(3)-invariant non-linear features to capture the complex non-linear patterns inherent in physical quantities.
@@
The non-linear neural mapping defined as the following is SO(3)-equivariant, satisfying the requirements specified in Eq. (4):
$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}}\quad\text{s.t.}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ (5)
where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, $s_{nonlin}(\cdot)$ represents non-linear neural modules, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
Proofs of the two theorems are included in Supplementary Information B.
Remark 1.
In Theorem 1, the term $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is equal to the square of the Frobenius norm of the matrix $\mathbf{Q}$, i.e., $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\|\mathbf{Q}\|_{F}^{2}=\sum_{m,n}|\mathbf{Q}_{mn}|^{2}$. The invariance of Frobenius norm under unitary transformations reflects the profound connections between modulus and invariance in physics.
Remark 2.
In Theorem 2, $g_{\text{nonlin}}(\cdot)$ can handle equivariant features of any degree and the returned feature $\mathbf{v}$ has the same degree $l$ as the input feature $\mathbf{f}$. $\mathbf{v}$ is endowed with non-linear expressiveness with the effect of $s_{nonlin}(\cdot)$, which can be any differentiable non-linear network module, including non-linear activation function as a part. The gradient operation $\frac{\partial z}{\partial\mathbf{f}}$ in $g_{\text{nonlin}}(\cdot)$ is used during the forward inference process to promote feature encoding, expanding the usage and significance of gradients within the deep learning paradigm.
Taking $z$ in Eq. (5) as the bridge connecting Theorem 1 and Theorem 2, we present a general method for learning non-linear representations that are equivariant to SO(3), as shown in Fig. 2. The core of our method can be abstractly referred to as TraceGrad. At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ introduced in Theorem 1 as a supervisory signal for learning SO(3)-invariant non-linear feature $z$. Given the attribute that invariance is preserved under non-linear operations, $z$ is encoded by $s_{nonlin}(\cdot)$ to obtain non-linear expressive capabilities. At the representation level, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as we prove. Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations to predict the SO(3)-equivariant regression target. For details of the method, please refer to Section 6.
@@
4 Results and Analyses
We apply our theory and method to the electronic-structure Hamiltonian prediction task, and collect results on eight benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), QH9-Stable (QS), and QH9-Dynamic (QD). The first six databases, consisting of crystal systems with elements like C, Mo, S, Bi, Te and Se, are from the DeepH benchmark series [6, 8]; while the last two databases, composed of molecular systems of C, H, O, N, F, are from the QH9 benchmark series [19] . These databases present diverse and complex challenges to the generality of a regression model. Regarding $MG$, $MM$, and QD, as their samples are prepared from a 300K environment, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction. For $BG$, $BB$, $BT$, and $BS$, their twisted samples with an interplay of SO(3)-equivariant effects and van der Waals (vdW) force variations bring significant generalization challenges, which are further exacerbated by the absence of any twisted samples in the training sets. Besides, $BB$, $BT$, and $BS$ exhibit strong spin-orbit coupling (SOC) effects, which further increase the complexity of Hamiltonian modeling. For QS and QD, as we select the most difficult subset split strategies provided in their paper, namely ’ood’ and ’mol’, respectively, yielding differences in atom numbers and trajectories between training and testing samples, challenges in scale and trajectory extrapolation become prominent. Detailed information of these databases, including their statistics and visualizations, can be found in the Supplementary Information C.
We use a comprehensive set of metrics to deeply evaluate the performance of electronic-structure Hamiltonian prediction models. On the databases from the DeepH benchmark series [6, 8], we adpot the metrics of Mean Absolute Error (MAE) between predicted and ground truth Hamiltonians introduced by Yin et al. [17], including $MAE^{H}_{all}$ for measuring average MAE of all samples and matrix elements, $MAE^{H}_{cha\_s}$ for measuring the MAE of challenging samples where the baseline model performs the worst, $MAE^{H}_{block}$ for measuring the MAE of different basic blocks in the Hamiltonian matrix, and $MAE^{H}_{cha\_b}$ for measuring the MAE on the most challenging Hamiltonian block where the baseline model shows the poorest performance (with the largest $MAE^{H}_{block}$). These metrics comprehensively reflect the accuracy performance, covering not only the average accuracy but also the accuracy on difficult samples and challenging blocks of the Hamiltonian matrices. On the two databases from the QH9 benchmark series, we adopt the metrics introduced by their original paper [19], including MAE of Hamiltonian matrices, which are further subdivided into $MAE^{H}_{all}$ for measuring average MAE, $MAE^{H}_{diag}$ for measuring MAE of Hamiltonian matrix formed by an atom with itself, and $MAE^{H}_{non\_diag}$ for measuring MAE of Hamiltonian matrices formed by different atoms; as well as the MAE ($MAE^{\epsilon}$) of occupied orbital energies $\epsilon$ induced by the predicted Hamiltonians and compared to the ground truth ones, and the cosine similarity ($Sim(\psi)$) between the electronic wavefunctions $\psi$ induced by the predicted and ground truth Hamiltonians. $\epsilon$ and $\psi$ are crucial downstream physical quantities for determining multiple properties of the atomic systems as well as their dynamics, highly reflecting the application values of the Hamiltonian regression model.
We compare experimental results from two setups: the first one is the experimental results of the baseline SO(3)-equivariant regression model [8, 7] for Hamiltonian prediction, and the second one is the experimental results of extending the architecture and pipeline of the baseline model through the proposed TraceGrad method, which incorporates non-linear expressiveness into the SO(3)-equivariant features of the baseline model with the gradient operations of SO(3)-invariant non-linear features learned by the trace targets. We choose DeepH-E3 [8] as the baseline model for databases from the DeepH benchmark series [6, 8]; and we choose QHNet [7] as the baseline model for databases from the QH9 benchmark series [19]. The results of DeepH-E3 to be compared are taken from the reproduction of Yin et al. [17]. The reproduction, based on DeepH-E3’s newest open-source project [^1] https://github.com/Xiaoxun-Gong/DeepH-E3, showed slightly better results than that in the original publication, and was evaluated more comprehensively across multiple metrics. The results of QHNet to be compared are taken from their original paper [19], and for the unification of MAE units, we convert the units of MAE from $10^{-6}$ Hartree ($E_{h}$) in the original paper to meV222 $27211.41E_{h}=$ 27211.4 meV. The results of our methods are the average from 10 independent repeated experiments. To ensure reproducibility, we used a fixed random seed for all random processes. As a result, the standard deviation of the Hamiltonian prediction MAE across repeated experiments does not exceed $0.007$ meV, which is negligible.
We list the results of DeepH-E3 and DeepH-E3+TraceGrad in Tables 1 and 2 for monolayer structure databases ($MG$, $MM$) and bilayer structure databases ($BG$, $BB$, $BT$, $BS$) from the DeepH benchmark series, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$. Regarding the metric of $MAE^{H}_{block}$ for every Hamiltonian block, due to its large data volume, we just present its values for representative datasets here in Fig. 3, and attach its values for all databases from the DeepH series in Supplementary Information E. From these results, we could find that the proposed TraceGrad method dramatically enhances the accuracy performance of the baseline method DeepH-E3, both on average and for challenging samples or blocks, both on the non-twisted samples and the twisted samples. Specifically, on the corresponding datasets, TraceGrad lowers down the $MAE^{H}_{all}$ and $MAE^{H}_{cha}$ of DeepH-E3 with relative ratios of up to $34\%$ and $35\%$, respectively, and achieving better results in the vast majority of basic blocks. Particularly, for the blocks where DeepH-E3 perform the worst, TraceGrad reduces the $MAE$ ($MAE^{H}_{cha\_b}$) by a maximum of $48\%$. The leading performance on the $MG$ and $MM$ databases prepared at 300K demonstrates the robustness of our method against thermal motion. The high accuracy on the $BB$, $BT$, and $BS$ databases, which have strong SOC effects, indicates our method’s strong capability to model such effects. The excellent performance on the $BG^{t}$, $BB^{t}$, $BT^{t}$, and $BS^{t}$ subsets showcases the method’s superior generalization to twisted structures, which are not present in the training data. The outstanding performance on such samples highlights the good potential for studying twist-related phenomena, given interlayer twists are hot research topics because they may bring new electrical and transport properties [20, 21, 22]. Additionally, the $BG^{t}$ and $BB^{t}$ testing subsets contain significantly larger atomic systems compared to the training set (see Supplementary Information C for statistics of their sizes), yet our method still excels on these subsets for all metrics, demonstrating its good scalability on the sizes of atomic systems it handles.
We list the results of QHNet and QHNet+TraceGrad in Tables 3 and 4, for the $QS$ and $QD$ databases, respectively, reporting the values of $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$. From these results, we can observe that the proposed TraceGrad method significantly enhances the accuracy performance of the baseline method QHNet across all of these metrics on both of the two databases. Specifically, TraceGrad lowers down the $MAE^{H}_{all}$, $MAE^{H}_{non\_diag}$, $MAE^{H}_{diag}$, $MAE^{\epsilon}$ and $Sim(\psi)$ of QHNet with relative ratios of up to 40%, 39%, 40%, 76%, and 17%, on the corresponding databases. The significant improvements on the $QS$ database, partitioned using the ’ood’ split strategy [19] without scales overlapping among the training, validation, and testing sets, once again demonstrate the method’s strong generalization capabilities across different scales. Meanwhile, the improvements on the $QD$ database under the ’mol’ strategy [19] partitioning the training, validation, and testing sets with samples from totally different thermal motion trajectories, highlight the method’s robustness in generalizing to new thermal motion sequences. Furthermore, the significant promotion on the prediction accuracy of $\epsilon$, i.e., occupied orbital energies which are crucial for determining electronic properties such as optical characteristics and conductivity in atomic systems, and $\psi$, i.e., the electronic wavefunctions essential for understanding electron distribution and interactions, underscore good application potential of our method for tasks like material design and quantum computing.
The experimental results above indicate that the proposed representation learning framework, while maintaining strict SO(3)-equivariance, significantly enhances the expressiveness of equivariant representations, thus excellently capturing the complexity of physical systems and substantially breakthroughs the SOTA accuracy on the eight benchmark databases. The consistent improvements brought by our framework to different baseline models (DeepH-E3 or QHNet) and in various challenging scenarios, such as thermal motions, bilayer twists, scale variations, and new trajectories, fully demonstrate the universality and significance of our proposed theory and method.
Table 1: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$, and $MAE^{H}_{cha\_b}$ metrics (meV) on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@ 4 Results and Analyses
Table 2: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$ and $MAE^{H}_{cha\_b}$ metrics (meV) on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$) and Bilayer Bi2Se3 ($BS$) databases, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
Figure 3: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for representative datasets, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. Lower value of $MAE^{H}_{block}$ corresponds to better accuracy. The $MAE^{H}_{block}$ for other datasets from DeepH series can be found in Supplementary Information E.
Table 3: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-Stable (QS) database using the ’ood’ split strategy [19]. $\downarrow$ means lower values of the metrics correspond to better accuracy, while $\uparrow$ means higher value of the metric corresponds to better performance. The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
Methods MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
@@
Table 4: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-Dynamic (QD) database using the ’mol’ split strategy [19]. $\downarrow$ means lower values of the metrics correspond to better accuracy, while $\uparrow$ means higher value of the metric corresponds to better performance. The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
Methods MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
$MAE^{H}_{all}$ $MAE^{H}_{diag}$ $MAE^{H}_{non\_diag}$ $MAE^{\epsilon}$
QHNet 4.733 11.347 4.182 264.483 0.792

### v4

@@ 4 Theory
The non-linear neural mapping $g_{nonlin}(\cdot)$ defined as the following is SO(3)-equivariant:
$\mathbf{v}=g_{nonlin}(\mathbf{f})=\frac{\partial z}{\partial\mathbf{f}},\quad\text{subject to}\quad z=s_{nonlin}(u),u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ (5)
where $\mathbf{f}$ is an input SO(3)-equivariant feature with degree $l$ in the direct-sum state, $\otimes$ denotes the direct-product operation of tensors, $\text{CGDecomp}(\cdot,0)$ refers to performing a Clebsch-Gordan decomposition of the tensor and returning the scalar component of degree 0, $s_{nonlin}(\cdot)$ represents arbitrary differentiable non-linear neural modules, and $\mathbf{v}$ is the outputted feature encoded by $g_{nonlin}(\cdot)$.
The proofs of the two theorems above are presented in Appendix B.
@@
5 Method
As shown in Fig. 1, taking $z$ in Eq. 5 serving as the bridge between Theorem 1 and Theorem 2, we propose a general method for learning non-linear representations that satisfy the SO(3)-equivariance property outlined in Eq. 4. The core of our method can be abstractly referred to as TraceGrad. At the label level, it incorporates the SO(3)-invariant trace quantity $tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ introduced in Theorem 1 as a supervisory signal for learning $z$, i.e., the SO(3)-invariant internal representations of $g_{\text{nonlin}}(\cdot)$ in Theorem 2. Given the attribute that invariance is preserved under non-linear operations, $z$ is encoded by $s_{nonlin}(\cdot)$ to obtain non-linear expressive capabilities. At the representation level, by taking the gradient of $z$ with respect to $\mathbf{f}$, the non-linear expressiveness of $z$ is transferred to the equivariant feature $\mathbf{v}$, while maintaining strict SO(3)-equivariance as we prove. Subsequently, merging $\mathbf{v}$ and $\mathbf{f}$, and applying $g_{\text{nonlin}}(\cdot)$ in a stacked manner can yield rich SO(3)-equivariant non-linear representations for inferring the SO(3)-equivariant regression target.
5.1 Encoding and Decoding Framework
Our encoding framework corresponds to a total of $K$ encoding modules, which are sequentially connected to form a deep encoding framework. For the $k$-th module ($1\leq k\leq K$), it first introduces an SO(3)-equivariant backbone encoder, e.g., the encoders from DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b) which is composed of recently developed equivariant operators (Thomas et al., 2018; Geiger & Smidt, 2022) like linear scaling, direct products, direct sums, Clebsch-Gordan decomposition, gated activation, equivariant normalization, and etc., to encode the physical system’s initial representations, such as spherical harmonics (Schrodinger, 1926), or representations passed from the previous neural layers, as equivariant feature $\textbf{f}^{(k)}$ in the current hidden layer. Next, we construct the feature $\textbf{v}^{(k)}=g_{\text{nonlin}}(\textbf{f}^{(k)})$, to achieve sufficient non-linear expressiveness while maintaining SO(3)-equivariance, where $g_{\text{nonlin}}(\cdot)$, $s_{nonlin}(\cdot)$ are the non-linear functions defined in Eq. 5. The function $s_{nonlin}(\cdot)$ can be implemented as any differentiable non-linear neural network module, such as feed-forward layers with non-linear activation functions like $SiLU$ and normalization operations like $Layernorm$.
For an expressive representation of a complex physical system, $\textbf{f}^{(k)}$ is usually not a feature of single degree but a direct-sum concatenation of series of components $\{\textbf{f}^{(k)l_{1}},\textbf{f}^{(k)l_{2}},\textbf{f}^{(k)l_{3}},...\}$ at multiple degrees, i.e., $L^{(k)}=\{l_{1},l_{2},l_{3},...\}$, where some components share the same degree while others differ. In this case, it becomes necessary to extend the decomposition operator, i.e., $\text{CGDecomp}(\cdot)$ in Eq. 5, to accommodate various components of degrees. Moreover, in the context of neural networks, introducing learnable parameters $W$ and more feature channels $C$ may improve the model capacity. Based on these considerations, when constructing the encoding module, the $\text{CGDecomp}(\cdot)$ operation in Eq. 5 can be expanded as $\text{CGDecomp}_{ext}(\cdot)$, and its outputs can be expanded as $\textbf{u}^{(k)}$, as shown in Eq. 6:
$\textbf{u}^{(k)}$ $=[u_{1}^{(k)},...,u_{c}^{(k)},...,u_{C}^{(k)}],$ (6)
$u_{c}^{(k)}$ $=\text{CGDecomp}_{\text{ext}}(\textbf{f}^{(k)}\otimes\textbf{f}^{(k)},0,W)=\sum_{l_{i},l_{j}\in L^{(k)},l_{i}=l_{j}}W^{c}_{ij}\cdot\text{CGDecomp}(\textbf{f}^{(k)l_{i}}\otimes\textbf{f}^{(k)l_{j}},0)$
where $W^{c}_{ij}$ represents learnable parameters, $u_{c}^{(k)}$ is the $c$ th channel of $\mathbf{u}^{(k)}$. $\text{CGDecomp}_{\text{ext}}(\cdot)$ adds learnable parameters to $\text{CGDecomp}(\cdot)$ and expands its output from a single channel to multiple channels. To further enhance the model capacity, we also expand the output of $s_{\text{nonlin}}(\cdot)$ to multiple channels as follows: $\mathbf{z}^{(k)}=[z_{1}^{(k)},\ldots,z_{c}^{(k)},\ldots,z_{C}^{(k)}]$. We define $\mathbf{v}_{c}^{(k)}=\frac{\partial z_{c}^{(k)}}{\partial\mathbf{f}^{(k)}}$, and construct the new features $\mathbf{v}^{(k)}$ by $\mathbf{v}^{(k)}=\sum_{c}\mathbf{v}_{c}^{(k)}$. It is evident that these extensions maintain the SO(3)-invariance of $\mathbf{u}^{(k)}$ and $\mathbf{z}^{(k)}$, as well as the SO(3)-equivariance of $\mathbf{v}^{(k)}$. $\mathbf{v}^{(k)}$ is combined with $\mathbf{f}^{(k)}$ in a residual manner like $\textbf{o}^{(k)}=\textbf{f}^{(k)}+\textbf{v}^{(k)}$ as the output of the $k$ th encoding module.
[…]
5.2 Training
The training loss function is shown as the following:
$\begin{split}\min_{\theta}loss&=loss_{Q}+\mu(loss_{Q},loss_{T})\cdot loss_{T},\\ loss_{Q}&=Error(\widehat{\mathbf{Q}},\mathbf{Q}^{*}),\,\,\,\,loss_{T}=Error(\widehat{\mathbf{T}},\mathbf{T}^{*})\end{split}$ (7)
where $\theta$ denotes all of the learnable parameters of our framework, $\widehat{\mathbf{Q}}$, $\widehat{\mathbf{T}}$ and $\mathbf{Q}^{*}$, $\mathbf{T}^{*}$ respectively denote the predictions and labels of $\mathbf{Q}$ and $\mathbf{T}$. In order to prevent the numerical disparity between the two loss terms and stabilize the training for both of the SO(3)-equivariant and SO(3)-invariant branches, we apply $\mu(loss_{Q},loss_{T})$, i.e., a coefficient to regularize the relative scale of the two loss terms:
where $No\_Grad(\cdot)$ denotes gradient discarding when calculating such coefficient, as this coefficient is only used for adjusting the relative scale between the two loss terms and should not itself be a source of training gradients, otherwise it would counteract the gradients from $loss_{T}$ in Eq. 7. All of the encoding and decoding modules are trained jointly by Eq. 7.
For the implementation details of our method, including the specific design of the network modules, parameter settings, and training specifics, please refer to Appendix D.
6 Experiments
6.1 Experimental Conditions
Table 1: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{cha\_s}$ and $MAE^{H}_{cha\_b}$ metrics (meV) on the Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$) and Bilayer Bi2Se3 ($BS$) databases, where the superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets.
@@ 6.1 Experimental Conditions
We apply our theory and method to the electronic-structure Hamiltonian prediction task, and collect results on eight benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), QH9-stable (QS), and QH9-dynamic (QD). The first six databases, consisting of periodic crystalline systems with elements like C, Mo, S, Bi, Te and Se, are from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023). The last two databases are from the QH9 benchmark series (Yu et al., 2023a), composed of molecular systems with elements like C, H, O, N and F. These databases present diverse and complex challenges to the generality of a regression model. Regarding $MG$, $MM$, and QD, as their samples are prepared from an temperature environment at three-hundred Kelvin, the thermal motions lead to complex non-rigid deformations, increasing the difficulty of Hamiltonian prediction. For $BG$, $BB$, $BT$, and $BS$, the twisted structures, with an interplay of SO(3)-equivariant effects and van der Waals (vdW) force variations bring significant generalization challenges, which are further exacerbated by the absence of any twisted samples in the training sets. Besides, $BB$, $BT$, and $BS$ exhibit strong spin-orbit coupling (SOC) effects, which further increase the complexity of Hamiltonian modeling. For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and testing sets, ensuring that the atom number of samples do not overlap across the three subsets. For the $QD$ database, the ’mol’ strategy provided by Yu et al. (2023a) is applied to split the training, validation, and testing sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets. The ’mol’ and ’ood’ strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories, respectively. Detailed statistic information of these databases can be found in the Appendix C. Implementation details of our method for experiments on these databases are presented in Appendix D.
Table 2: Experimental results measured by the $MAE^{H}_{all}$, $MAE^{H}_{diag}$, $MAE^{H}_{non\_diag}$, $MAE^{\epsilon}$, and $Sim(\psi)$ metrics on the QH9-stable (QS) and QH9-dynamic (QD) databases respectively using ’ood’ and ’mol’ split strategies (Yu et al., 2023a). $\downarrow$ means lower values correspond to better accuracy, while $\uparrow$ means higher values correspond to better performance. The units of MAE metrics are meV, while $Sim(\psi)$ is the cosine similarity which is dimensionless.
Methods QS
MAE ()($\downarrow$)( ↓ ) $Sim(\psi)$ ()($\uparrow$)( ↑ )
@@
Methods QD
QHNet (Baseline) 4.733 11.347 4.182 264.483 0.792